OpenStudy (anonymous):

Perimeter is 76in and the length is 23in the width of the rectangle is

OpenStudy (lgbasallote):

perimeter = 2(length) + 2(width) 2 width = perimmeter - 2(length) width = (P - 2L)/2 where L = length; P = perimeter

OpenStudy (anonymous):

so whats the answer

OpenStudy (lgbasallote):

substitute :P haha

OpenStudy (p0sitr0n):

answer is 15

OpenStudy (anonymous):

w=15 ans